Circular No. 6958

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998cu IN IC 1525 W. D. Li, M. Modjaz, R. R. Treffers, and A. V. Filippenko, University

of California at Berkeley, report their discovery of an apparent supernova during the course of the Lick Observatory Supernova Search (cf. IAUC 6627) with the 0.8-m Katzman Automatic Imaging Telescope (KAIT), as part of a cooperative program with the Beijing Astronomical Observatory (BAO) Supernova Survey (cf. IAUC 6379). SN 1998cu was found and confirmed in unfiltered CCD images obtained on June 29.4 (mag ∼ 18.4) and 30.4 UT (mag ∼ 18.1). The new object is located at α = 23^{h}59^{m}12^{s}.95, δ = +46^{o}52′31′′.8 (equinox 2000.0), which is ∼ 29′′.2 west and 50′′.3 south of the nucleus of IC 1525. BAO images (limiting mag ∼ 18.5) obtained on May 29 and June 17 do not show a star at the position of the supernova; a KAIT image of IC 1525 taken on 1997 Dec. 26 (limiting mag ∼ 19.0) also shows nothing at this position.

XTE J2123–058 T. Takeshima, Goddard Space Flight Center (GSFC) and Universities

Space Research Association; and T. E. Strohmayer, GSFC, on behalf of the RXTE Science Operations Center, write: “We report an improved position for this transient (IAUC 6955) of α = 21^{h}23^{m}16^{s}, δ = −5^{o}47′.5 (equinox 2000.0), with an uncertainty of 1′, based on scans across the source with the Proportional Counter Array on RXTE on June 29.04 UT. This is consistent with the position of the candidate of an optical counterpart reported on IAUC 6957. Two weak bursts were seen during the pointed observations on June 27.98 and 29.70. Preliminary spectral analysis suggests that they are type-I (thermonuclear) bursts; thus the system is likely a low-mass xray binary with a neutron star. The He II emission reported by Tomsick et al. (IAUC 6957) also supports this hypothesis. The observed steady fluxes (2–10 keV) on June 27.09, 29.04, and 29.70 were 65, 67, and 56 mCrab, respectively. The observed burst peak fluxes above the steady flux level (2–10 keV) were 50 and 88 mCrab.”

COMET C/1998 H1 (STONEHOUSE) Visual m_{1} estimates (L = reflector): May 14.92 UT, 11.1 (G. Comello,

Roden, The Netherlands, 0.30-m L); 23.12, 11.5 (J. Bortle, Stormville, NY, 0.41-m L); June 2.89, 12.5 (M. Lehký, Hradec Králové, Czech Republic, 0.42-m L); 29.89, 13.8 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m L).

1998 June 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green